# Hydrodynamic models of Reheating

### v1

@@ I Introduction
As we shall show below, consistency requirements at the macroscopic level alone are not enough to single out an unique hydrodynamical model. To fill the gaps, we shall assume hydrodynamics is obtained from a more fundamental description, where the fluid excitations may be described as quasiparticles with a definite mass, obeying a kinetic equation of motion [52]. For tractability, we employ the relaxation-time approximation for the collision integral in kinetic theory [53, 45].
@@
In the present model, the degrees of freedom include the usual temperature and velocity (with vanishing chemical potential, as appropriate in the absence of a matter–antimatter asymmetry) as well as two additional tensor fields. The first is a second-rank tensor associated with the viscous part of the energy–momentum tensor. The second is a third-rank tensor, required to allow tensor fluctuations to propagate with a definite speed [54, 57, 58].
@@ Appendix A Derivation of the hydrodynamical equations
In this Appendix we derive the hydrodynamical equations of the model (1). We consider a relativistic causal fluid within the framework of second-order theories, where dissipative dynamics are treated as genuine degrees of freedom [85]. Such theories can be derived from a kinetic-theory formulation, in which the Boltzmann equation takes the form
@@
We include an external force given by $F_{\mu}=-\tfrac{1}{2}\partial_{\mu}M^{2}$, with $M$ the fluid mass, representing an effective coupling between the fluid and the condensate. We allow for a field-dependent fluid mass $M=M(\phi)$, which may also exhibit a temperature dependence.
We consider a massive fluid with vanishing chemical potential whose 1pdf is parametrized as
@@
$\int Dp\,S^{\alpha\beta}_{\mu\nu}\frac{p^{\mu}p\nu}{(-u_{\lambda}p^{\lambda})}\left[p^{\tau}f_{;\tau}-I_{col}\right]$ $=-F^{\tau}\int Dp\,S^{\alpha\beta}_{\mu\nu}\frac{p^{\mu}p^{\nu}}{(-u_{\lambda}p^{\lambda})}\frac{\partial}{\partial p^{\tau}}f$
@@
where $Dp$ is the invariant measure in momentum space (10).
@@
Thus, the first-order dynamical equations take the form (5) with the definitions (6) and (7). This scheme allows us to write the equations for the background and the tensor sector of the fluid fluctuations as
@@
In this work we only consider tensor perturbations of the fluid and the Minkowski metric, the latter describing the gravitational waves as $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$. Finally, Eqs. (73) can be recast as Eqs. (20) and (31), which describe the background dynamics and the tensor perturbations, respectively.

### v2

@@ I Introduction
As we shall show below, consistency requirements at the macroscopic level alone are not enough to single out an unique hydrodynamical model. To fill the gaps, we shall assume hydrodynamics is obtained from a more fundamental description, where the fluid excitations may be described as quasiparticles with a definite mass, obeying a kinetic equation of motion [13, 56, 54, 12, 14]. For tractability, we employ the relaxation-time approximation for the collision integral in kinetic theory [9, 21].
@@
In the present model, the degrees of freedom include the usual temperature and velocity (with vanishing chemical potential, as appropriate in the absence of a matter–antimatter asymmetry) as well as two additional tensor fields. The first is a second-rank tensor associated with the viscous part of the energy–momentum tensor. The second is a third-rank tensor, required to allow tensor fluctuations to propagate with a definite speed [52, 11, 85, 18, 26, 25, 84].
@@ Appendix A Derivation of the hydrodynamical equations
In this Appendix we derive the hydrodynamical equations of the model (1). We consider a relativistic causal fluid, where dissipative dynamics are treated by including auxiliary degrees of freedom [89].
Such theories can be derived as the low frequency, long wavelength limit of a quantum field theory, where the field excitations are described as quasi-particles with a temperature dependent mass [13, 55, 56, 54, 12, 14].
The gas of quasiparticles obeys a Boltzmann equation
@@
The external force is $F_{\mu}=-\tfrac{1}{2}\partial_{\mu}M^{2}$, where $M$ the quasi-particle mass, which depends both on the temperature of the quasi-particles and the amplitude of the condensate. Because of particle-antiparticle symmetry we assume the gas has vanishing chemical potential.
We shall further reduce the kinetic description to a hydrodynamic one by applying the method of moments [65]. We first choose a preferred set of functions $c_{A}$, and then compute their moments
$C^{\mu}_{A}=\int Dp\;p^{\mu}c_{A}f,$ (66)
where the invariant measure $Dp$ is defined in eq. (10). The $C^{\mu}_{A}$ are regarded as the fundamental observables of the theory. Because the energy-momentum tensor plays a special role, it is natural to choose the first function $c_{0}=p_{\nu}$. Then, further moments are added in order to obtain a causal and stable theory [97, 27, 52, 11, 26, 84, 101, 53].
Once the $c_{A}$ functions are chosen, to obtain a close system of equations the 1pdf is approximated by the maximum entropy distribution function, subject to equations (66) as so many constraints
$f\approx e^{\sum_{A}\lambda_{A}c_{A}}$ (67)
whereby a set of Lagrange multipliers $\lambda_{A}$ are introduced. Given our choice of $c_{0}$ it is natural to choose $\lambda_{0}=\beta_{\nu}$, which is further identified with the inverse temperature four vector. The Lagrange multipliers become the dynamical degrees of freedom of the theory, since the $C^{\mu}_{A}$ moments are functions of them.
We obtain the dynamical equations by taking moments of the kinetic equation
$\int Dp\;c_{A}\left[p^{\nu}\nabla_{\nu}f-I_{\text{col}}+F^{\nu}\frac{\partial f}{\partial p^{\nu}}\right]=0.$ (68)
These equations imply positive entropy production [19].
Equations (68) define the divergence of the moments (66) as linear combinations of derivatives of the Lagrange multipliers, multiplied by suitable transport coefficients. These coefficients may be derived by fitting the hydrodynamic response functions to the low frequency, long wavelength limit of the underlying quantum field theory [85, 53].
Since we assume that the collision term takes the relaxation-time-approximation form (see Eq. (77) below), characterized by a single relaxation time $\tau$, all transport coefficients are proportional to it. However, the theory can be straightforwardly generalized to allow for multiple relaxation times [40, 51]. We discuss the relation between the relaxation time and the shear-viscosity transport coefficient in Appendix (C); analogous relations for the other transport coefficients can be derived following a procedure similar to that of [26, 53].
In this work our choice of the preferred moments shall be
$c_{0}$ $=p^{\mu}$ (69)
$c_{1}$ $=S^{\alpha\beta}_{\mu\nu}\frac{p^{\mu}p^{\nu}}{(-u_{\lambda}p^{\lambda})}$
and
$c_{2}=S^{\alpha\beta\gamma}_{\mu\nu\rho}\frac{p^{\mu}p^{\nu}p^{\rho}}{(-u_{\lambda}p^{\lambda})^{2}}$ (70)
where the projectors $S^{\alpha\beta}_{\mu\nu}$ and $S^{\alpha\beta\gamma}_{\mu\nu\rho}$ are defined in eq. (2). Thereby we must include the Lagrange multipliers
$\lambda_{0}$ $=\beta_{\mu}$ (71)
$\lambda_{1}$ $=\zeta_{\mu\nu}$
$\lambda_{2}$ $=\xi_{\mu\nu\rho}$
where $\zeta_{\mu\nu}$ and $\xi_{\mu\nu\rho}$ are totally symmetric, transverse and traceless. The powers of $(-u_{\lambda}p^{\lambda})$ appearing in $c_{1}$ and $c_{2}$ are introduced because, in test cases where kinetic theory and hydrodynamic calculations can be performed simultaneously and directly compared, they yield significantly improved agreement between the two formalisms [19, 85, 18]. As a consequence, our final equations differ from those presented in [84, 26, 101, 53]. Nevertheless, the different sets of equations are easily seen to describe the same physics.
In summary, we parametrize the 1pdf as
@@
$\int Dp\,S^{\alpha\beta}_{\mu\nu}\frac{p^{\mu}p^{\nu}}{(-u_{\lambda}p^{\lambda})}\left[p^{\tau}f_{;\tau}-I_{col}\right]$ $=-F^{\tau}\int Dp\,S^{\alpha\beta}_{\mu\nu}\frac{p^{\mu}p^{\nu}}{(-u_{\lambda}p^{\lambda})}\frac{\partial}{\partial p^{\tau}}f$
@@
where $Dp$ is the invariant measure in momentum space (10). Note that, in principle, when integrating these equations to obtain eq. (1), one must also take into account derivatives acting on the integration measure. These contributions are given by
$\int\frac{d^{4}p}{(2\pi)^{3}}\,c_{A}f\left[p^{\nu}\nabla_{\nu}+F^{\nu}\frac{\partial}{\partial p^{\nu}}\right]\delta(p^{2}+M^{2})=\int\frac{d^{4}p}{(2\pi)^{3}}\,c_{A}p^{\nu}f\left(M^{2}_{;\nu}+2F_{\nu}\right)\delta^{\prime}(p^{2}+M^{2}),$ (74)
where $\delta^{\prime}$ denotes the derivative of the Dirac delta function with respect to its argument. These terms vanish identically due to the specific choice of the external force.
@@
Thus, up to first order, the tensors in (75) that appear in the conservation laws (1) are
$A^{\mu\nu\rho}$ $=T_{1}^{\mu\nu\rho}+2C^{5}_{2}(2)u^{\mu}\zeta^{\nu\rho}+6C^{6}_{3}(3)\xi^{\mu\nu\rho}$ (80)
$A^{\mu\nu\rho\sigma}$ $=T_{2}^{\mu\nu\rho\sigma}+2C^{6}_{3}(2)\left\{u^{\mu}u^{\nu}\zeta^{\rho\sigma}\right\}+2C^{6}_{3}(3)\left\{\Delta^{\mu\nu}\zeta^{\rho\sigma}\right\}+6C^{7}_{4}(3)\left\{u^{\mu}\xi^{\nu\rho\sigma}\right\}$
$A^{\mu\nu\rho\sigma\lambda}$ $=T_{3}^{\mu\nu\rho\sigma\lambda}+2C^{7}_{4}(2)\{u^{\mu}u^{\nu}u^{\rho}\zeta^{\sigma\lambda}\}+2C^{7}_{4}(3)\{u^{\mu}\Delta^{\nu\rho}\zeta^{\sigma\lambda}\}$
$+6C^{8}_{5}(3)\{u^{\mu}u^{\nu}\xi^{\rho\sigma\lambda}\}+6C^{8}_{5}(4)\{\Delta^{\mu\nu}\xi^{\rho\sigma\lambda}\}$
$I^{\mu\nu}$ $=\frac{2}{\tau}\,C^{4}_{1}(2)\,\zeta^{\mu\nu}$
$I^{\mu\nu\rho}$ $=\frac{6}{\tau}\,C^{6}_{3}(3)\,\xi^{\mu\nu\rho}$
$B^{\mu}$ $=T_{1}^{\mu}$
$B^{\mu\nu}$ $=T_{2}^{\mu\nu}+2C^{4}_{3}(2)\zeta^{\mu\nu}$
$B^{\mu\nu\rho}$ $=T_{3}^{\mu\nu\rho}+2C^{5}_{4}(2)\{u^{\mu}\zeta^{\nu\rho}\}+6C^{6}_{5}(3)\xi^{\mu\nu\rho}$
where the tensors appearing inside the braces are understood to be completely symmetrized. Here we defined
$T_{m}^{\mu_{1}...\mu_{n}}=\int Dp\,\frac{p^{\mu_{1}}\ldots p^{\mu_{n}}}{(-u_{\lambda}p^{\lambda})^{m}}f_{0}.$ (81)
These tensors, after integration, can be written in terms of the fluid four-velocity and the coefficients $C^{n}_{m}(k)=a^{2k}_{n-m-2k}/(2k-1)!!$, as in (6). The functions $a^{2k}_{n-m-2k}$ were defined in eq. (8). Then, the first-order dynamical equations take the form (5). This scheme allows us to write the equations for the background and the tensor sector of the fluid fluctuations as
@@
In this work we only consider tensor perturbations of the fluid and the Minkowski metric, the latter describing the gravitational waves as $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$. Finally, the background evolution can be reduced to the single equation (20), while the tensor dynamics, after including the metric perturbation, becomes eqs. (31).